OpenStudy (anonymous):

Prove that i^98+w^99=0 . . where i=√(-1) and w is the complex cube root of unity.

OpenStudy (anonymous):

w^99-1=0 w^99=1 1=1

OpenStudy (anonymous):

w^99-1=0 how??

OpenStudy (anonymous):

what happened to i

OpenStudy (phi):

i^0 = 1 i^1 = i i^2 = -1 I^3 = -i (from i*i^2= i*-1) i^4 = 1 (i*i^3= i*-i= -(i*i)= -(-1) = +1. or i^2*i^2= -1*-1= 1) i^5 = i there is a pattern here... notice we can rewrite as i^98= i^96 * i^2 i^96= (i^4)^24 * i^2 i^4= 1 and 1^24 is 1 i^2= -1 1*-1= -1 for W^99, we are told W^3= 1 rewrite W^99= (W^3)^33 = 1^33= 1

OpenStudy (phi):

I used rules for exponents \[ (x^a)^b = x^{ab}\] and \[ x^{a+b}=x^a\cdot x^b \]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how it is coming -1

OpenStudy (phi):

If you don't like using exponents, you could write out i^98 = i*i*i*i....*i 98 times put parens around each group of 4 i's: i^98= (i*i*i*i) * (i*i*i*i) * ... (i*i*i*i) * (i*i*i*i) * i*i you will get 24 groups of i*i*i*i each = 1 and i*i left over: 1*1*1... *1 * -1= -1

OpenStudy (anonymous):

(i^4)^24=-1 how

OpenStudy (phi):

(i^4)^24 = 1 notice (i^4)^24 = i^96 we want i^98

OpenStudy (phi):

maybe the big numbers are confusing... if you had i^6= i*i*i*i * i*i (that is what i to the 6th power means) but we know i*i*i*i= 1 (see above) so i^6= 1* i*i but i*i= -1 (by definition) so i^6= 1* -1 = -1 now if we had i^10 it would be i^10 = i*i*i*i * i*i*i*i * i*i but this is i^10 = 1 * 1 * -1 i^10= -1 each time we multiply by i^4, we are multiplying by 1

OpenStudy (anonymous):

pretty confusing

OpenStudy (anonymous):

but got it

OpenStudy (phi):

think about it, it should make sense.

OpenStudy (anonymous):

w^99=1 how

OpenStudy (anonymous):

Hint : i^4 =1 w^3 =1

OpenStudy (anonymous):

i know this

OpenStudy (anonymous):

just need to ask w^99 =1 how?

OpenStudy (phi):

you know w^3= 1 (w is the cube root of 1) so think like this w*w*w * w*w*w * w*w*w * .... in groups of 3 you will have exactly 33 groups each w*w*w = 1 and 1 times itself 33 times is still 1

OpenStudy (anonymous):

i^98 = (i^4)^24 * i^2

OpenStudy (anonymous):

w^99 = (w^3)^33 = 1 1^2 = -1 so -1+1 =0

OpenStudy (anonymous):

thanks alot